OpenStudy (anonymous):

Need help with an undetermined coefficient problem. Will fan if all steps are shown! Problem inside.

OpenStudy (anonymous):

\[y'''+y'=2+sinx\]

OpenStudy (anonymous):

homogeneous: y'''+y'=0 k^3+k=0 <--characteristic equation k(k^2+1)=0 k=0 or k=i or k=-i y=e^(0t) or y=e^(it) or y=e^(-it) y=c or y=c e^(it) or y=ce^(-it) particular: sin(x)=(ie^(-ix)-ie^(ix))/2 yp=c4 x e^(ix) + c5 x e^(-ix) + c6 x <-- x's are added in so that the particular differs from the homogeneous solutions solve for c4,c5,c6 then y=yp+c1+c2 e^(it)+c3 e^(-it)

OpenStudy (anonymous):

Thanks. Question: I thought for the particular you use the f(x)

OpenStudy (anonymous):

you do use the f(x) for the particular except when the f(x) is equal to the homogeneous. in that case, you need to multiply by x; otherwise, the particular will be 0 since the homogeneous is 0. so for f(x)=2+sin(x) the first term, 2, needs to be multiplied by x (c6 x) otherwise y'''+y' for y=c*2 will be 0 since y=c is a homogeneous solution similarly, for the e^(it) component of sin(x), y'''+y' for y=c e^(it) will be 0 since y=c e^(it) is a homogeneous solution so we need to use y=c*x*e^(it) for the particular, so that y'''+y' is not equal to 0

OpenStudy (anonymous):

also on an unrelated note, you can express all the complex exponentials in terms of sin and cosine if you prefer these over imaginary numbers

OpenStudy (anonymous):

yeah I did, can I show you what I have done so far?

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

\[y'''+y'=2+sinx\] \[y_c=c_1+c_2cosx+c_3+sinx\]

OpenStudy (anonymous):

\[y_p=Asinx+B\] \[y'_p=Acosx\] \[y''_p=-Asinx\] \[y'''_p=-Acosx\]

OpenStudy (anonymous):

is this right?

OpenStudy (anonymous):

close, however, the particular cannot be equal to the complementary if y=Asinx, then y'''+y' will equal 0 (since sinx is a complementary solution) to fix this, multiply by x: yp=A x sin(x) B has the same problem yp=A x sin(x)+Bx

OpenStudy (anonymous):

the way you have written it, yp is a subset of yc

OpenStudy (anonymous):

Ohhhhhhh

OpenStudy (anonymous):

So is there a case where I don't have to tack on an X.

OpenStudy (anonymous):

yup, if the the complementary solution did not have a c3 sin(x) term, you could write yp=Asin(x)+Bx similarly, if the complementary did not have a c1 term, yp=Axsin(x)+B would work and if the complementary had neither a c1 term not a c3 sin(x) term, then yp=Asin(x)+B would be fine

OpenStudy (anonymous):

*nor

OpenStudy (anonymous):

But since your y_p is derived from your F(x), aren't you essentially copying it every time?

OpenStudy (anonymous):

f(x) is independent of the characteristic equation though the characteristic equation is what determines the complementary solution if the complementary solution and F(x) share terms, then you need to multiply these shared terms by x

OpenStudy (anonymous):

but yes, you are copying F(x) each time and potentially multiplying terms by x or taking derivatives

OpenStudy (anonymous):

Oh I think I get it now...So if r values provide an equation similar to F(x), I must tack on an X to avoid cancellation in future steps, right?

OpenStudy (anonymous):

I wish I could give you another star lol...you don't how long I was confused about this lol.

OpenStudy (anonymous):

yup :) also, http://tutorial.math.lamar.edu/Classes/DE/UndeterminedCoefficients.aspx explains the method in detail and provides lots of examples

OpenStudy (anonymous):

glad i was able to help :)

OpenStudy (anonymous):

yeah my friend was telling me about that site. Thanks again. You may see me posting a lot tonight though...test tomorrow XD anyway appreciate it!